Solve the equation. 4 -(x+6) = 4x - 7
Answer by TimothyLamb(4379)

---
4 - (x + 6) = 4x - 7
4 - x - 6 = 4x - 7
4 - x - 6 + 7 = 4x - 7 + 7
4 - x + 1 = 4x
4 + 1 - x = 4x
5 - x = 4x
5 - x + x = 4x + x
5 = 5x
5/5 = 5x/5
1 = x
---
answer:
x = 1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php